OpenStudy (anonymous):

Find the area of a parallelogram if a base and corresponding altitude have the indicated lengths. Base 1 1/2 feet, altitude 6 inches.

OpenStudy (anonymous):

You can do this! :P Remember, area = base*height

OpenStudy (anonymous):

You have different values here, so try converting them both to the same value. Ie both to feet or both to inches. Remember, 1 foot = 12 inches

OpenStudy (anonymous):

So would it be 3/4 in. ?

OpenStudy (anonymous):

1.5 feet = 18 inches. Area = base*height Area = 18*6 = 108inches squared

OpenStudy (anonymous):

or.. 6 inches = 0.5 feet Area = base*height Area = 1.5*0.5 = 0.75 feet squared.

OpenStudy (anonymous):

Okay thanks i get it now